# Defining Cost Function of Steganography with Large Language Models

### v2

@@ Introduction
In this paper, unlike the work in [10] that both the carrier and the secret information are presented explicitly within the prompt, we employ LLMs to generate computer programs capable of facilitating steganographic tasks. We aim to enable large models to discover novel steganographic cost functions in the form of computer programs. In detail, a two-stage strategy combining LLM-guided program synthesis with evolutionary search is proposed in this paper. The first stage generates a certain number of candidate cost functions. The second stage determines the present optimal cost function(s). These two stages are performed by an iterative fashion so that the best cost function is determined at the last iteration. Experimental results show that the proposed method enables LLMs to design new cost functions of steganography that significantly outperform existing works in terms of resisting steganalysis tools, which verifies the superiority and feasibility. This work provides a novel insight into the algorithmic design of steganography and has the potential to guide further research.
@@
To automate the design of steganographic algorithms, we introduce a novel two-stage framework that integrates LLMs with an evolutionary (code) search mechanism. The proposed framework evolves over multiple generations, each comprising two important stages including cost function generation and cost function evaluation. As shown in Figure 1, Stage 1 employs prompt-guided LLMs to iteratively generate and refine code for cost functions, which are subjected to rapid preliminary evaluation. Stage 2 conducts a more rigorous and accurate assessment of the cost functions produced in Stage 1 and provides targeted feedback for refinement. Through repeated evolution across generations, this framework accumulates a diverse set of candidate functions and progressively identifies superior cost functions under robust evaluation. Different from conventional, manually designed heuristic methods, the proposed framework enables automated exploration of expansive search spaces and continuously adjusts its search trajectory through feedback, thereby substantially enhancing the efficiency and creativity of algorithmic design. We show more details about the two stages in the following subsections.
@@
The objective of the Stage 1 is to efficiently identify and preliminarily filter a set of candidate cost functions, ensuring both the scalability and search efficiency of the framework. To achieve this objective, an iterative evolution-based search strategy driven by LLM is employed. With structured prompts, the LLM iteratively refines existing cost functions to produce multiple improved variants. These variants then undergo rapid evaluation, and those that execute successfully and yield valid scores are incorporated into the program database as candidate cost functions for steganography. All cost functions are in the form of computer programs. We describe the workflow of Stage 1 in the following.
@@
In other words, the initial program is present in each sub-database. In fact, each $\text{PD}_{i}$ consists of multiple clusters of programs, which will be clarified later. Obviously, each $\text{PD}_{i}$ now consists of only one cluster, i.e., $\{p_{\text{raw}}\}$. A reference program, denoted by $p_{\text{ref}}$, is thereafter sampled from $S_{\text{P}}$ as follows: First, an index $1\leq i\leq n$ is randomly generated; Then, $p_{\text{ref}}$ is sampled from some cluster of $\text{PD}_{i}$ based on the temperature-scaled softmax function used to determine the selection probability of each cluster. Clearly, suppose that $\text{PD}_{i}=\{\text{C}_{i,1},\text{C}_{i,2},...,\text{C}_{i,k_{i}}\}$, where $k_{i}>0$ is the total number of clusters in $\text{PD}_{i}$. Each $\text{C}_{i,j}$ is associated with a score $s_{i,j}$. Here, a larger $s_{i,j}$ implies that the cost functions in $\text{C}_{i,j}$ have better performance resisting against the steganalysis tools. We will provide more information about $s_{i,j}$ later. In this way, the probability of sampling $p_{\text{ref}}$ from $\text{C}_{i,j}$ is given by
@@
which suggests a shorter program as $p_{\text{ref}}$ as longer programs are more prone to errors in terms of no matter computer programming or cost function definition.
@@
The above operations will be performed in an iterative manner, resulting in an increasing number of programs in the database. To improve efficiency, after performing the above operations multiple times, we remove ‘bad’ programs from the database. First, all the sub-databases are marked as ‘good’ or ‘bad’ based on their highest cluster scores. That is, for each sub-database, if its highest cluster score is no less than a threshold, it is marked as ‘good’, otherwise it is marked as ‘bad’. Then, each ‘bad’ sub-database is replaced with a new sub-database that only contains the highest-score cluster of a randomly selected ‘good’ sub-database. Thus, all sub-databases become ‘good’, i.e., most programs are ‘good’. For implementation, it may be not easy to specify the value of the threshold. An alternative strategy is to set the percentage of ‘bad’ sub-databases to be replaced, e.g., 50%, which has been applied in our experiments. In conclusion, it is always free for us te design the strategy updating the database so as to improve efficiency.
@@
Remark: After multiple times of evolutionary, programs that achieve good performance in both preliminary and accurate evaluation, are marked as potential functions. Obviously, these potential functions are all in the program database. They are then evaluated across different embedding rates. For each of these candidate cost functions, we train steganalyzers using different embedding rates and perform accurate evaluation for each embedding rate. A candidate is only accepted as a superior cost function if it outperforms the initial program across all embedding rates. This multi-rate validation is essential because it enforces cross-rate generalization and prevents overfitting to a single rate, ensuring that the final promoted functions reflect genuine and robust security improvements. If there are multiple superior cost functions, any one can be selected as the final cost function generated by the LLM. One may also design a new strategy to select one of them.
@@
The number of sub-databases is set as two, i.e., $|\text{S}_{\text{P}}|=n=2$. The temperature parameter $T$ in Eq. (10) is set as 0.1 initially, and then gradually decreases as the number of iterations increases. In each iteration, the framework samples only one reference function (i.e., $r=1$), which is then combined with a fixed natural language prompt. The resulting prompt is therefore fed into the LLM four times to generate four new evolved candidate functions, i.e., $n_{p}=4$. After an initial burn-in period of ten iterations, the process enters Stage 2 for accurate evaluation, subsequently re-entering Stage 2 every five iterations, ensuring frequent assessment to effectively guide the search. To avoid long-term stagnation caused by low-quality programs, the program database is refreshed every four hours (here, we use the running time to determine whether to update the program database) by removing ‘bad’ programs based on the preliminary evaluation scores of corresponding functions.
@@
$\textbf{K}_{\text{e}}^{(0)}=\textbf{d}_{1}\otimes\textbf{d}_{0}=\frac{1}{4}\begin{bmatrix}1&1\\ -1&-1\end{bmatrix}$ (16)
$\textbf{K}_{\text{e}}^{(1)}=\textbf{d}_{0}\otimes\textbf{d}_{1}=\frac{1}{4}\begin{bmatrix}1&-1\\ 1&-1\end{bmatrix}$ (17)
$\textbf{K}_{\text{e}}^{(2)}=\textbf{d}_{1}\otimes\textbf{d}_{1}=\frac{1}{4}\begin{bmatrix}1&-1\\ -1&1\end{bmatrix}$ (18)
$\textbf{K}_{\text{e}}^{(3)}=\textbf{d}_{0}\otimes\textbf{d}_{0}=\frac{1}{4}\begin{bmatrix}1&1\\ 1&1\end{bmatrix}$ (19)
@@
$\rho_{i,j}=\left\{\begin{matrix}+\infty&\text{if}~\rho_{i,j}^{\prime}\leq\theta,\\ \rho_{i,j}^{\prime}&\text{if}~\rho_{i,j}^{\prime}>\theta,\end{matrix}\right$. (24)
@@
$\rho=\frac{1}{\xi}\otimes\textbf{L}_{2}.$ (31)
@@
$\rho_{\text{e}}=\frac{1}{\xi}\otimes\textbf{G}_{\sigma}.$ (32)
@@
$\xi^{(k)}=\left(\frac{1}{|\textbf{R}^{(k)}|+\epsilon}\right)\odot\left|\textbf{K}^{(k)}\right|,~k=0,1,2.$ (34)
The final embedding cost is then given by
@@
$\rho_{i,j}^{(-1)}=\left\{\begin{matrix}+\infty&\text{if}~x_{i,j}<\tau,\\ \rho_{i,j}&\text{if}~x_{i,j}\geq\tau.\end{matrix}\right$. (39)
@@
We report the performance of the evolved algorithms against steganalysis. The performance is evaluated across various embedding rates ranging from 0.1 bpp to 0.4 bpp with a step of 0.1. The detection error rate in Eq. (13) is adopted as the indicator. The experimental setting is the same as the process of evolutionary, e.g., SRNet is used as the steganalyzer, and the ratio between training set, validation set and test set is 5:1:4. The experimental results shown in Table 1 indicate that all evolved algorithms consistently outperform their original counterparts, exhibiting enhanced security across every embedding strength and improved resistance to steganalysis. To quantify this improvement, we determine the (relative) performance gain, defined as $\Delta^{\uparrow}=(P_{\text{E}}^{\star}-P_{\text{E}}^{\circ})/P_{\text{E}}^{\circ}$, where $P_{\text{E}}^{\star}$ and $P_{\text{E}}^{\circ}$ correspond to the evolved algorithm and raw version.

### v3

@@ Introduction
Unlike the work in [10] where both the carrier and the secret information are presented explicitly within the prompt, we employ LLMs to generate computer programs capable of facilitating steganographic tasks. We aim to enable large models to discover novel steganographic cost functions in the form of computer programs. In detail, a two-stage strategy combining LLM-guided program synthesis with evolutionary search is proposed in this paper. The first stage generates a certain number of candidate cost functions. The second stage determines the present optimal cost function(s). These two stages are performed by an iterative fashion so that the best cost function is determined at the last iteration. Experimental results show that the proposed method enables LLMs to design new cost functions of steganography that significantly outperform existing works in terms of resisting steganalysis tools, which verifies the superiority and feasibility. This work provides a novel insight into the algorithmic design of steganography and has the potential to guide further research.
@@
To automate the design of steganographic algorithms, we introduce a novel two-stage framework that integrates LLMs with an evolutionary (code) search mechanism. The proposed framework evolves over multiple generations, each comprising two important stages, namely, cost function generation and cost function evaluation. As shown in Figure 1, Stage 1 employs prompt-guided LLMs to iteratively generate and refine code for cost functions, which are subjected to rapid preliminary evaluation. Stage 2 conducts a more rigorous and accurate assessment of the cost functions produced in Stage 1 and provides targeted feedback for refinement. Through repeated evolution across generations, this framework accumulates a diverse set of candidate functions and progressively identifies superior cost functions under robust evaluation. Different from conventional, manually designed heuristic methods, the proposed framework enables automated exploration of expansive search spaces and continuously adjusts its search trajectory through feedback, thereby substantially enhancing the efficiency and creativity of algorithmic design. We show more details about the two stages in the following subsections.
@@
The objective of Stage 1 is to efficiently identify and preliminarily filter a set of candidate cost functions, ensuring both the scalability and search efficiency of the framework. To achieve this objective, an iterative evolution-based search strategy driven by LLM is employed. With structured prompts, the LLM iteratively refines existing cost functions to produce multiple improved variants. These variants then undergo rapid evaluation, and those that execute successfully and yield valid scores are incorporated into the program database as candidate cost functions for steganography. All cost functions are in the form of computer programs. We describe the workflow of Stage 1 in the following.
@@
In other words, the initial program is present in each sub-database. In fact, each $\text{PD}_{i}$ consists of multiple clusters of programs, which will be clarified later. Obviously, each $\text{PD}_{i}$ now consists of only one cluster, i.e., $\{p_{\text{raw}}\}$. A reference program, denoted by $p_{\text{ref}}$, is thereafter sampled from $S_{\text{P}}$ as follows: First, an index $1\leq i\leq n$ is randomly generated; Then, $p_{\text{ref}}$ is sampled from some cluster of $\text{PD}_{i}$ based on the temperature-scaled softmax function used to determine the selection probability of each cluster. Clearly, suppose that $\text{PD}_{i}=\{\text{C}_{i,1},\text{C}_{i,2},...,\text{C}_{i,k_{i}}\}$, where $k_{i}>0$ is the total number of clusters in $\text{PD}_{i}$. Each $\text{C}_{i,j}$ is associated with a score $s_{i,j}$. Here, a larger $s_{i,j}$ implies that the cost functions in $\text{C}_{i,j}$ have better performance resisting against the steganalysis tools. We will provide more information about $s_{i,j}$ later. In this way, the probability of sampling $\text{C}_{i,j}$ from $\text{PD}_{i}$ is given by
@@
which suggests a shorter program as $p_{\text{ref}}$ as longer programs are more prone to errors in terms of both computer programming and cost function definition.
@@
The above operations will be performed in an iterative manner, resulting in an increasing number of programs in the database. To improve efficiency, after performing the above operations multiple times, we remove ‘bad’ programs from the database. First, all the sub-databases are marked as ‘good’ or ‘bad’ based on their highest cluster scores. That is, for each sub-database, if its highest cluster score is no less than a threshold, it is marked as ‘good’, otherwise it is marked as ‘bad’. Then, each ‘bad’ sub-database is replaced with a new sub-database that only contains the highest-score cluster of a randomly selected ‘good’ sub-database. Thus, all sub-databases become ‘good’, i.e., most programs are ‘good’. For implementation, it may be not easy to specify the value of the threshold. An alternative strategy is to set the percentage of ‘bad’ sub-databases to be replaced, e.g., 50%, which has been applied in our experiments. In conclusion, it is always free for us to design the strategy updating the database so as to improve efficiency.
@@
Remark: After multiple rounds of evolution, programs that achieve good performance in both preliminary and accurate evaluation, are marked as potential functions. Obviously, these potential functions are all in the program database. They are then evaluated across different embedding rates. For each of these candidate cost functions, we train steganalyzers using different embedding rates and perform accurate evaluation for each embedding rate. A candidate is only accepted as a superior cost function if it outperforms the initial program across all embedding rates. This multi-rate validation is essential because it enforces cross-rate generalization and prevents overfitting to a single rate, ensuring that the final promoted functions reflect genuine and robust security improvements. If there are multiple superior cost functions, any one can be selected as the final cost function generated by the LLM. One may also design a new strategy to select one of them.
@@
The number of sub-databases is set to two, i.e., $|\text{S}_{\text{P}}|=n=2$. The temperature parameter $T$ in Eq. (10) is set to 0.1 initially, and then gradually decreases as the number of iterations increases. In each iteration, the framework samples only one reference function (i.e., $r=1$), which is then combined with a fixed natural language prompt. The resulting prompt is therefore fed into the LLM four times to generate four new evolved candidate functions, i.e., $n_{p}=4$. After an initial burn-in period of ten iterations, the process enters Stage 2 for accurate evaluation, subsequently re-entering Stage 2 every five iterations, ensuring frequent assessment to effectively guide the search. To avoid long-term stagnation caused by low-quality programs, the program database is refreshed every four hours (here, we use the running time to determine whether to update the program database) by removing ‘bad’ programs based on the preliminary evaluation scores of corresponding functions.
@@
$\textbf{K}_{\text{e}}^{(0)}=\textbf{d}_{1}\cdot\textbf{d}_{0}^{\text{T}}=\frac{1}{4}\begin{bmatrix}1&1\\ -1&-1\end{bmatrix}$ (16)
$\textbf{K}_{\text{e}}^{(1)}=\textbf{d}_{0}\cdot\textbf{d}_{1}^{\text{T}}=\frac{1}{4}\begin{bmatrix}1&-1\\ 1&-1\end{bmatrix}$ (17)
$\textbf{K}_{\text{e}}^{(2)}=\textbf{d}_{1}\cdot\textbf{d}_{1}^{\text{T}}=\frac{1}{4}\begin{bmatrix}1&-1\\ -1&1\end{bmatrix}$ (18)
$\textbf{K}_{\text{e}}^{(3)}=\textbf{d}_{0}\cdot\textbf{d}_{0}^{\text{T}}=\frac{1}{4}\begin{bmatrix}1&1\\ 1&1\end{bmatrix}$ (19)
@@
$\rho_{i,j}=\left\{\begin{matrix}+\infty&\text{if}~\rho_{i,j}^{\prime}<\theta,\\ \rho_{i,j}^{\prime}&\text{if}~\rho_{i,j}^{\prime}\geq\theta,\end{matrix}\right$. (24)
@@
$\rho=\frac{1}{\xi+\epsilon}\otimes\textbf{L}_{2},$ (31)
where $\epsilon$ can be set to, e.g., $10^{-10}$.
@@
$\rho_{\text{e}}=\frac{1}{\xi+\epsilon}\otimes\textbf{G}_{\sigma}.$ (32)
@@
$\xi^{(k)}=\left(\frac{1}{|\textbf{R}^{(k)}|+\epsilon}\right)\odot\left|\textbf{K}^{(k)}\right|,~k=0,1,2,$ (34)
where $\epsilon$ is set to 1. The final embedding cost is then given by
@@
$\rho_{i,j}^{(+1)}=\left\{\begin{matrix}+\infty&\text{if}~x_{i,j}\geq 255-\tau,\\ \rho_{i,j}&\text{if}~x_{i,j}<255-\tau,\end{matrix}\right$. (38)
@@
$\rho_{i,j}^{(-1)}=\left\{\begin{matrix}+\infty&\text{if}~x_{i,j}\leq\tau,\\ \rho_{i,j}&\text{if}~x_{i,j}>\tau.\end{matrix}\right$. (39)
@@
We report the performance of the evolved algorithms against steganalysis. The performance is evaluated across various embedding rates ranging from 0.1 bpp to 0.4 bpp with a step of 0.1. The detection error rate in Eq. (13) is adopted as the indicator. The experimental setting is the same as the evolutionary process, e.g., SRNet is used as the steganalyzer, and the ratio between training set, validation set and test set is 5:1:4. The experimental results shown in Table 1 indicate that all evolved algorithms consistently outperform their original counterparts, exhibiting enhanced security across every embedding strength and improved resistance to steganalysis. To quantify this improvement, we determine the (relative) performance gain, defined as $\Delta^{\uparrow}=(P_{\text{E}}^{\star}-P_{\text{E}}^{\circ})/P_{\text{E}}^{\circ}$, where $P_{\text{E}}^{\star}$ and $P_{\text{E}}^{\circ}$ correspond to the evolved algorithm and raw version.